### v2

# Towards More Realistic Extraction Attacks: An Adversarial Perspective

Abstract
Language models are prone to memorizing parts of their training data which makes them vulnerable to extraction attacks. Existing research often examines isolated setups–such as evaluating extraction risks from a single model or with a fixed prompt design. However, a real-world adversary could access models across various sizes and checkpoints, as well as exploit prompt sensitivity, resulting in a considerably larger attack surface than previously studied. In this paper, we revisit extraction attacks from an adversarial perspective, focusing on how to leverage the brittleness of language models and the multi-faceted access to the underlying data. We find significant churn in extraction trends, i.e., even unintuitive changes to the prompt, or targeting smaller models and earlier checkpoints, can extract distinct information. By combining information from multiple attacks, our adversary is able to increase the extraction risks by up to 2×2$\times$2 ×. Furthermore, even with mitigation strategies like data deduplication, we find the same escalation of extraction risks against a real-world adversary. We conclude with a set of case studies, including detecting pre-training data, copyright violations, and extracting personally identifiable information, showing how our more realistic adversary can outperform existing adversaries in the literature.Equal contribution [^2] Code released at https://github.com/EQUAL-Mila/llm_extraction_eval
1 Introduction
Large language models (LLMs) have grown considerably in size (Meta AI, 2024; Zhao et al., 2023), and have become integral to a wide range of tasks such as knowledge retrieval, question answering, code generation, machine translation, etc.
To complement this growing scale, LLMs are often trained on large amounts of data (Penedo et al., 2024; Soboleva et al., 2023; Gao et al., 2020; Raffel et al., 2020) that may include private, unlicensed or copyrighted information, especially if directly scraped from the web. As LLMs are prone to memorizing the data they’ve been trained on, they can be prompted to expose sensitive contexts - making it easier for an adversary to extract information in a black-box setting. Naturally, a question arises, how big is the risk imposed due to memorization?
Extraction attacks offer an empirical framework to quantify the information leakage in the presence of an adversary. The most commonly studied extraction attack is discoverable memorization (Carlini et al., 2023; Kassem et al., 2024), where the model is prompted with a portion of a sentence from the training data to extract the rest, thus enabling the adversary to perform targeted attacks.
Current extraction attacks study memorization trends in LLMs across isolated settings like model sizes, generation hyperparameters and learning dynamics (Carlini et al., 2021). While effective, they underestimate the risk posed due to a multi-faceted access to the underlying data in the current LLM ecosystem. For instance, we show that an adversary can exploit the sensitivity of LLMs to prompt structure, length and content, to amplify the information gained. The current accessibility to frequently updated model sizes (Meta AI, 2024); checkpoints (Biderman et al., 2023b; Groeneveld et al., 2024) and a large array of model families such as Llama (Meta AI, 2024), Gemini  (Team et al., 2023), and Falcon  (Almazrouei et al., 2023), can also create higher extraction risks.
In this paper, we study a more realistic scenario and explore the actual risks posed by extraction attacks. More specifically, we ask:
  1. 1.
    Can adversaries exploit prompt sensitivity? We find that extraction attacks are sensitive to the prompt design, extracting over $20\%$ more data with even minor, unintuitive changes to the prompt (§5.1). Thus, an adversary, given the opportunity to prompt the model multiple times, can extract more data than previously observed.
  2. 2.
    Does access to multiple checkpoints increase extraction risk? An adversary with access to multiple model checkpoints over time or sizes gains broader access to the underlying dataset. We find such an adversary can increase the extraction rates up to 1.5×1.5$\times$1.5 ×, significantly heightening the risk of information leakage (§5.2).
  3. 3.
    Is data deduplication effective in reducing the extraction risks? We find that data deduplication does reduce the extraction risks, in line with the existing literature (Carlini et al., 2023). However, adversaries can still exploit the prompt structure and multiple checkpoints to extract more information (§6.3). Thus, our concerns about a powerful real-world adversary persist despite deduplication.
  4. 4.
    How are downstream applications affected by the presence of such an adversary? We performed three separate case studies and found that our more realistic adversary improves the $p$-value of dataset inference up to 2×2$\times$2 ×7.1), the extraction of copyright violations by up to $20\%$7.2), and the extraction rate of personally identifiable information (PIIs) by 1.5×1.5$\times$1.5 ×7.3).
2 Background and Related Work
In this section, we introduce relevant background on extraction attacks in LLMs, followed by an overview of related work on prompt sensitivity and training dynamics in LLMs. Finally, we describe the term churn as it applies in our context.
Extraction Attacks in LLMs. Unintended memorization in LLMs can make it prone to information leakage (Tirumala et al., 2022; Carlini et al., 2019; Mattern et al., 2023; Carlini et al., 2022), particularly through extraction attacks (Birch et al., 2023; Carlini et al., 2021, 2023; Nasr et al., 2023). These attacks allow adversaries to extract training data from the model, raising concerns of leaking sensitive information (Birch et al., 2023). Extraction attacks have gained significant attention in recent years, studied under two primary frameworks: Discoverable Memorization (Carlini et al., 2023; Kassem et al., 2024; Liu et al., 2023b; Biderman et al., 2023a; Tirumala et al., 2022; Huang et al., 2022), where the adversary attempts to extract targeted information, and Extractable Memorization Nasr et al. (2023); Kandpal et al. (2022); Qi et al. (2024), where the adversary attempts to extract any information about the data.
We add to the growing body of research on targeted extraction attacks by highlighting the lack of a realistic adversary in the literature. We show the existence of a stronger real-world adversary capable of combining information from various attacks, thereby defining a composite form of discoverable memorization (§3). Schwarzschild et al. (2024) also redefines discoverable memorization, using prompt optimization and adversarial compression ratio (ACR) to quantify memorization as information compression. In contrast, rather than relying on optimization, our focus is instead on exploiting the multi-faceted access to LLM training data.
Prompt Sensitivity in LLMs. LLMs are shown to be sensitive to changes in their prompts, leading to fluctuations in their performance (Sclar et al., 2024; Liu et al., 2023a). This sensitivity persists across varying model sizes and through fine-tuning and other downstream modifications (Salinas and Morstatter, 2024; Zhu et al., 2023). The sensitivity of prompts can also be misused, and adversarial modifications to prompts can trigger the model to act in unintended ways Rossi et al. (2024); Liu et al. (2024); Hubinger et al. (2024); Liu et al. (2024). While several overarching trends studying the impact of prompt design on extraction attacks are present in the literature (Carlini et al., 2023; Kassem et al., 2024; Qi et al., 2024; Tirumala et al., 2022), these trends are often evaluated in isolation. Motivated by the composability of privacy leakage (McSherry, 2009), we argue that an adversary capable of repeated prompting can combine these trends. We show that such an adversary can extract more information about the training data than previously reported in the literature (§5.1).
Training Dynamics of LLMs. Several recent works have studied the training dynamics of LLMs over time (Tirumala et al., 2022; Liu et al., 2021; Xia et al., 2023). In the context of memorization, recent work by Biderman et al. (2023a) explored the impact of model size and intermediate checkpoints on the dynamics of memorization, revealing a considerable variance in memorized data over time and size. The practice of releasing models in various sizes and regularly updating them over time can thus increase the attack surface for the underlying dataset. In our work, we study how adversaries can exploit access to multiple checkpoints of a model to extract more data (§5.2).
Figure 1: Composability in LLMs. In the real world, an adversary has multi-faceted access to a dataset by (a) exploiting prompt sensitivity, and (b) accessing multiple checkpoints trained on the same data.
Churn. The instability of model predictions under updates has gained significant attention in recent years, studied under the umbrella of churn (Milani Fard et al., 2016; Cotter et al., 2019; Bahri and Jiang, 2021; Anil et al., 2018; Jiang et al., 2021; Adam et al., 2023; Watson-Daniels et al., 2024). Churn quantifies the inconsistency in predictions between a system pre-update vs post-update, by measuring the fraction of examples whose predictions diverge (Milani Fard et al., 2016). The term churn is traditionally used in the literature to describe such regressive trends in model predictions, we extend its use by highlighting similar regressive trends and instability of extraction attacks under changing prompts and models. Thus, churn occurs when information is extractable with weaker setups like shorter prompts, smaller models, or earlier checkpoints, but not with the stronger setup.
3 Re-evaluating Adversarial Strengths
The adversary is central to our work. We begin by defining its capabilities, arguing that existing work has underestimated the strength of real-world adversaries. To ensure broad applicability, we assume gray-box access to the target model, i.e., the adversary can only access the generation output and probabilities from the model. Consequently, they cannot access model weights, gradients, or even control the generation hyperparameters, which reflects the typical level of accessibility for most commercial LLMs. Despite these constraints, we will demonstrate that an adversary in the current LLM ecosystem possesses far greater power than what has been recognized in existing literature.
3.1 Adversary Capabilities
Composability (or self-composability) of privacy leakage (McSherry, 2009) suggests that when an adversary gains access to multiple outputs from algorithms on the same underlying dataset—whether through multiple queries from the same algorithm or queries across multiple algorithms—the risk of information leakage grows. Consequently, an adversary with multiple points of access is significantly stronger than one with only a single point of access. In the current landscape of LLMs, such access is not only unsurprising but also easily obtainable (as illustrated in Figure 1). Specifically, we consider two forms of multi-faceted access:
Exploiting Prompt Sensitivity LLMs are highly sensitive to their input, including its structure, content, and even the presence of noisy text within the prompt (Sclar et al., 2024; Liu et al., 2023a; Salinas and Morstatter, 2024; Zhu et al., 2023). While existing studies have focused on improving the prompts for stronger attacks, the nuance of prompt sensitivity in LLMs often defies intuitive expectations. For instance, while longer prompts are known to increase the success of extraction attacks Carlini et al. (2023), our work demonstrates that even shorter prompts can at times exploit vulnerabilities that longer prompts overlook (§5.1).
Given the widespread use of LLMs through both chat interfaces and API calls, restricting model access is not realistic. While most commercial LLMs do have rate limits, they are quite high to be of practical concern. For example, even at the lowest tier subscription of \$$\$5$, ChatGPT has a 500 query per minute (qpm) rate limit for GPT4 and 3500 qpm for GPT3.5333qpm stats and subscription rate as of September 2024.. Thus, an adversary can prompt millions of generations in just one day, making it easier to exploit structural changes in prompts.
Multiple Checkpoints. LLMs are typically deployed in various sizes to cater different needs for accuracy and efficiency among users. However, due to the stochastic nature of their training and the impact of scaling, different model sizes might memorize unique portions of the underlying dataset (Biderman et al., 2023a). Consequently, an adversary with access to multiple model sizes can effectively aggregate extracted information rather than limiting it to a single model (§5.2).
Similarly, deployed LLMs undergo regular updates driven by new data, better learning techniques, evolving security measures, and novel functionalities. The stochastic training process means that data resilient to attacks at a certain time step may become vulnerable in subsequent model updates, or vice-versa (Biderman et al., 2023a). Such fluctuations can enable adversaries to exploit multiple checkpoints over time, potentially extracting more information than from a static model (§5.2).
More broadly, access to multiple models sharing common training data increases the attack surface, and in turn, creates stronger adversaries. This level of access is not unprecedented, and several companies in the current LLM ecosystem release multiple versions of their models and even update them periodically. For example, there are 8 different major versions of the ChatGPT models and more than 10 major versions of the Llama models currently available, while these models are also known to get regular minor updates accessible using update dates OpenAI (2024); Chen et al. (2023). Thus, access to multiple models trained on the same data, as has become commonplace, can significantly increase the risks of information leakage.
3.2 Combining Extraction Attacks
We argued for the heightened risk posed by multifaceted access to LLMs, either through repeated prompting or multiple model checkpoints. Before discussing our empirical study, we first quantify the risks associated with this stronger adversary. We argue that when an adversary gains such extensive access, any successful extraction of information—even if achieved once—renders that specific information vulnerable to the adversary.
Formally, adapting the definition of discoverable memorization from Nasr et al. (2023), we propose:
Definition 3.1 (Composite Discoverable Memorization).
For a set of $k$ models $\mathbb{G}=\{Gen_{i}|i\in[1\dots k]\}$, a set of $r$ prompt modifiers $\mathbb{F}=\{F_{j}|j=[1\dots r]\}$, and an example $[{\mathbf{p}}\parallel{\mathbf{x}}]$ from the training set $\mathbb{X}$, we say ${\mathbf{x}}$ is composite discoverably memorized if $\exists\;Gen_{i}\in\mathbb{G}\;\textrm{and}\;F_{j}\in\mathbb{F}\;\textrm{s.t.}\;Gen_{i}(F_{j}(\mathbf{p}))=\mathbf{x}$.
$CDM(\mathbb{G},\mathbb{F},$ ${\mathbf{p}}\parallel{\mathbf{x}})$
$=\max_{Gen_{i}\in\mathbb{G},F_{j}\in\mathbb{F}}\mathbbm{1}_{Gen_{i}(F_{j}(\mathbf{p}))=\mathbf{x}}$
Prompt modifiers are defined as functions $F_{j}:\mathcal{W}^{*}\rightarrow\mathcal{W}^{*}$ that take a prompt as input and return a modified version of this prompt as output. Here, $\mathcal{W}$ represents a finite set of all tokens in the training data i.e $\mathcal{W}=\{w_{1},w_{2},\dots,w_{n}\}$ with $w_{i}$ representing individual tokens, and $\mathcal{W}^{*}$ represents the Kleene star operation over $\mathcal{W}$, i.e., a set of all finite length sequences (strings) of tokens in $\mathcal{W}$.
Figure 2: Choosing prompts (pre $100k$ steps) and checkpoints (post $100k$ steps) for evaluation of Pythia.
Extraction attacks are often evaluated in the literature using a verbatim match  (Carlini et al., 2021, 2023; Nasr et al., 2023; Huang et al., 2022), i.e., the generated text must match the original text perfectly. However, this rigid metric does not take into account the noise in LLM generations, and several recent works have turned to approximate matching to quantify extraction risks for LLMs (Qi et al., 2024; Kassem et al., 2024; Liu et al., 2023b; Ippolito et al., 2022). Thus, we also extend our definition of composite extraction attacks to the approximate matching setup:
Definition 3.2 (Approximate Composite Discoverable Memorization).
For a set of $k$ models $\mathbb{G}=\{Gen_{i}|i\in[1\dots k]\}$, a set of $r$ prompt modifiers $\mathbb{F}=\{F_{j}|j=[1\dots r]\}$, a similarity metric $S$, a similarity threshold $\delta$, and an example $[{\mathbf{p}}\parallel{\mathbf{x}}]$ from the training set $\mathbb{X}$, we say ${\mathbf{x}}$ is approximate composite discoverably memorized if $\exists\;Gen_{i}\in\mathbb{G}\;\textrm{and}\;F_{j}\in\mathbb{F}\;\textrm{s.t.}\;S\left(Gen_{i}\left(F_{j}\left(\mathbf{p}\right)\right),\mathbf{x}\right)\geq\delta$.
$ACDM(\mathbb{G},\mathbb{F},$ $S,\delta,{\mathbf{p}}\parallel{\mathbf{x}})$
$=\max_{Gen_{i}\in\mathbb{G},F_{j}\in\mathbb{F}}\mathbbm{1}_{S\left(Gen_{i}\left(F_{j}\left(\mathbf{p}\right)\right),\mathbf{x}\right)\geq\delta}$
Here, $S$ is a similarity metric defined as a function $S:(\mathcal{W}^{*}\times\mathcal{W}^{*})\rightarrow[0,1]$ that takes as input two strings $a,b\in\mathcal{W}^{*}$, and returns a score between 00 and $1$ to represent the similarity between the two input strings, and $\delta$ is a threshold that controls the degree of approximate matching.
4 Experimental Setup
In this section, we outline our central experiment setup, to set the stage for our empirical study. Note that details about the setup for the case studies (§7) are delegated to their respective sections.
4.1 Models and Dataset
We use the Pythia suite (Biderman et al., 2023b) and OLMo models (Groeneveld et al., 2024) for all our experiments. We primarily focus on the Pythia suite, which contains decoder-only language models with the same architecture as EleutherAI’s GPT-Neo (Black et al., 2022), albeit different training, across various model sizes and with open source access to the complete training data and the intermediate checkpoints. Pythia models were trained using GPT-NeoX library (Andonian et al., 2023) on the Pile dataset (Gao et al., 2020) and have not undergone any form of instruction-tuning. The standard version of Pythia was trained over a single epoch of the Pile dataset, i.e., $\approx 143k$ steps with a batch size of $1024$, while the deduplicated version of Pythia was trained over $\approx 1.5$ epochs of the deduplicated Pile dataset, maintaining the same number of training steps as the standard version.
Pythia suite of models was developed with an emphasis on facilitating open-source investigation into the training dynamics of LLMs. As such, they offer access to (a) models of various sizes (we use model sizes: 1b, 1.4b, 2.8b, 6.9b, and 12b), (b) intermediate model checkpoints during training (a total of $154$ checkpoints, with $144$ of them equally spaced, i.e., at every $1k$ training steps), and (c) the complete training data order, which is the same for all model sizes. This level of accessibility and control over the training setup allows us to simulate the real-world availability of models across various sizes and with updating checkpoints over time.
To show the generalizability of our results, we also perform some additional experiments with OLMo models, another set of decoder-only language models. OLMo models were trained on the Dolma dataset Soldaini et al. (2024) and have also not undergone any form of instruction-tuning. These models also offer access to (a) intermediate model checkpoints during training, and (b) open-source access to the complete training data order.
5050505020020020020035035035035050050050050000%$1$%$2$%$3$%$4$%Composite Ext. Rate Prompt Length Extraction Rate (a) Extraction rate over Prompt Length
1b1.4b2.8b6.9b12b00%$0.5$%$1$%$1.5$%Composite Ext. Rate Model Size p (b) Extraction rate over Model Size p
100k110k120k130k140k00%$0.5$%$1$%$1.5$%$2$%Composite Ext. Rate Training Step (c) Extraction rate over Training Steps
10030050000%$1$%$2$%$3$% Prompt Length Extraction Rate (d) Extraction rate over Masking and Removal
BaselineMask ($5\%$)Mask ($10\%$)Remove ($5\%$)Remove ($10\%$)Composite Prefix00%$0.5$%$1$%Prefix g Extraction Rate (e) Extraction rate over Prefixes g No PrefixNumbersAlphanumericComposite
Figure 3: Extraction rates under prompt sensitivity and across multiple models for Pythia. (a) Increasing prompt length results in better extraction rates, with the composite extraction rate better than even at prompt length $500$. (b, c) We see similar trends for increasing model size and training steps, respectively. Specifically, we see the largest impact of the composite extraction rate across training steps, with the extraction rate increased 1.5×1.5$\times$1.5 × compared to any single checkpoint. (d) Randomly masking or removing tokens from the prompt severely hurts the extraction rate, highlighting the importance of prompt structure. (e) Adding a random prefix can also contribute to minor improvements in the composite extraction rate.
4.2 Evaluation Methodology
We now describe our approach to the design and evaluation of extraction attacks. Similar to Carlini et al. (2023), we sample a representative portion of the dataset for analyzing the performance of our extraction attacks. More specifically, we uniformly sample $100,000$ sequences from the first $100k$ steps (batches) of the training data for Pythia. This sampling strategy is important because we choose model checkpoints for evaluation starting at step $100k$, which ensures that every sentence evaluated for memorization has been seen by each checkpoint under consideration, as illustrated in Figure 2. We use the same approach for OLMo, with the training step $300k$ being the cut-off point.
Each sequence sampled is exactly $2049$ tokens. For our analysis, we employ a consistent method of partitioning each sequence into a prompt and completion at the midpoint, i.e., $1024$ tokens. Formally, for a sentence $s_{1:2049}$, prompt length $l_{p}$, and completion length $l_{x}$, the example $[{\mathbf{p}}\parallel{\mathbf{x}}]$ is defined as ${\mathbf{p}}=s_{1024-l_{p}:1024}$ and ${\mathbf{x}}=s_{1024:1024+l_{x}}$. This partitioning allows us to systematically vary the prompt length and design while comparing the same completion, and vice-versa.
For the Pythia suite, unless otherwise specified, we use a prompt length of $l_{p}=50$, a completion length of $l_{x}=50$, the Pythia-6.9b model, and the 140k training step checkpoint, evaluating the extraction attacks using verbatim match. We use the same default setup for OLMo, with the OLMo-7b model, and the 500k training step checkpoint.
5 Churn in Extraction Trends
Churn (Milani Fard et al., 2016), as previously introduced in §2, refers to regressive variance for individual extracted information despite an overall improvement in the extraction rates. For instance, although using a longer prompt is often associated with stronger extraction rates (Carlini et al., 2023; Biderman et al., 2023a), we observe trends that exhibit churn, i.e., certain information is instead extractable only with shorter prompts but not with longer prompts. These non-monotonic and locally regressive trends of certain sentences (i.e., churn) can be exploited by an adversary with multifaceted access to the data to execute a composite extraction attack. We study the factors that may lead to churn such as (a) prompt sensitivity, and (b) access to models of varying sizes and training checkpoints.
5.1 Prompt Sensitivity
We start by examining prompt sensitivity, focusing on how trends in prompt design can lead to churn.
Prompt Length. Prompt length is a commonly studied parameter in extraction attacks, and it has been shown that longer prompts lead to better extraction (Carlini et al., 2023). This is intuitive, as conditioning the model with more text from training would increase the likelihood of extraction. We will now show that the composite extraction rate (Definition 3.1) across varying prompt lengths exceeds the extraction rate at even the largest prompt length. As illustrated in Figure 3(a), the extraction rate increases with longer prompts. However, the composite extraction rate is noticeably higher than any single prompt length, including the longest at 500 tokens. This suggests that certain information extractable with shorter prompts remains elusive even with the longest prompt. Consequently, an adversary can exploit this churn across the prompt length to extract more information. We see similar trends for OLMo in Figure 4.
Prompt Length Training Steps 00%$0.5$%$1$%$1.5$%$2$%$2.5$% Extraction Rate
Prompt Length $50$Training Step $300k$Prompt Length $500$Training Step $500k$Composite All
Figure 4: Composite extraction attack results across 10 prompt lengths (same as Pythia) and 11 training steps (equidistant between $300k$ and $500k$), compared against isolated setups, for OLMo.
Prompt Structure. Next, we explore the structure of prompts to identify what makes a prompt potent and where churn can emerge. We introduce noise into the prompts by masking and removing random tokens; results are collected in Figure 3(d). Despite introducing only a small amount of noise, we observe a significant drop in extraction rates. This indicates that the contiguous prompt from the training data is crucial for extracting information, and any disruption inside this prompt can significantly hurt its capabilities. Yet, we do see minute churn in extraction trends, which further highlights how an adversary can exploit repeated prompting to extract more information, even with seemingly unintuitive changes like masking or removing random tokens.
We next add noise as a prefix in the form of random numeric and alphanumeric strings; results are collected in Figure 3(e). Interestingly, the performance degradation with a noisy prefix is less severe than with noise within the prompt. More importantly, we observed a higher composite extraction rate. This suggests that adding a noisy prefix can also help extract unique information that was previously inaccessible, and further highlights how an adversary can exploit repeated prompting.
Note that the churn in our prompt design trends highlights the increased extraction risks without access to new information. For instance, if an adversary has access to the prompt of length 500 tokens, they can expand the attack surface and thereby the extraction rate simply by removing parts of the prompt, adding noise, etc.—without needing any additional knowledge. One might argue that as the number of prompt variations increases, every sentence could become extractable. However, that is not true; not all sentences are extractable. Yin et al. (2024) showed that knowledge not present in an LLM will not be extractable even after prompt optimization, while Schwarzschild et al. (2024) also showed similar trends when attempting to extract a given completion. Consequently, prompting an LLM to regurgitate certain sentences, even with various prompt modifications, demonstrates a genuine extraction risk and underscores the extent of memorization in LLMs (Carlini et al., 2021).
5.2 Multiple Checkpoints
Model Size. The model size has long been known to influence learning trends, and our results in Figure 3(b) reflect this phenomenon. We find that larger models tend to memorize more information, which makes them more vulnerable to extraction attacks. However, our results also indicate that the composite extraction rate is higher than the extraction rate of any single model, highlighting the churn present in these trends. Biderman et al. (2023a) also conducted an empirical study on the overlap between memorized data across model sizes and found that up to $10\%$ of the data memorized by smaller models is not memorized by larger models. Combining our insights with existing literature, it’s clear that releasing models in different sizes increases the extraction risks.
(a) Composite Extraction Rates Across Multiple Axes
1.19%1.53%1.85%1.71%2.11%2.50%2.03%2.49%2.90%Prompt LengthModel Size1.4b+2.8b+6.9b100+200+300
1.19%1.71%2.03%1.41%1.96%2.29%1.52%2.08%2.42%Training StepsPrompt Length100+200+300100k+120k+140k
1.19%1.41%1.52%1.53%1.76%1.87%1.85%2.12%2.27%Model SizeTraining Steps100k+120k+140k1.4b+2.8b+6.9b
PilePile-Deduped00%$1$%$2$%$3$% Dataset Extraction Rate (b) gComposite Attacks and Deduplicationg
[…]
Figure 5: Towards more realistic extraction rates by combining various churn trends and with approximate matching. (a) Combining two axes at a time, we see a monotonically increasing trend in extraction rates as we gain more points of access to the underlying dataset, highlighting the growing power of the adversary. (b) Combining multiple axes of attack results in a significant increase in extraction rate for both standard and deduplicated setups, with the composite extraction rate for the deduplicated setup the same as the highest single setting rate for the standard setup. (c) Various similarity metrics have distinct trends as we decrease the threshold value and allow for looser approximations, thus the choice is context-driven.
Model Updates. We also analyze model updates over time using intermediate checkpoints in Figure 3(c), where we observe the most significant churn in our study. Unsurprisingly, attacking models at later stages of training is more successful, as seen in the literature (Tirumala et al., 2022; Biderman et al., 2023a; Jagielski et al., 2023). But remarkably, the churn here is significantly powerful and by exploiting composability across intermediate checkpoints, an adversary can increase their extraction rate by more than 1.5×1.5$\times$1.5 ×. We also see similar results for OLMo in Figure 4. This underscores the impact of stochasticity in model training on extraction attacks and reveals that regular model updates, typically considered beneficial in the current LLM ecosystem, create a powerful adversary.
6 Towards Realistic Extraction Attacks
With a better understanding of the trends across various setups, we now evaluate a more realistic measure of leakage in extraction attacks, by investigating (a) composability in churn, (b) challenges in evaluation, and (c) effects of deduplication.
6.1 Combining Multiple Axes of Churn
In the previous section, we saw how churn can impact individual axes of variability, such as prompt sensitivity, model size, and intermediate checkpoints. However, a real-world adversary can take advantage of all these factors simultaneously, thus significantly increasing their extraction rates. We start by analyzing two axes at a time, as shown in Figure 5(a). For all pairs of variability, the overall composite extraction rate (bottom right) is 23×2-3$\times$2 - 3 × higher than the base setup (top left) and 1.52×1.5-2$\times$1.5 - 2 × higher than the composite extraction rates along one axis (top right and bottom left). Furthermore, when all three axes are combined, depicted in Figure 5(b), the extraction rates grow even higher, albeit with diminishing gains. Thus, we show that a real-world adversary can extract far more training data than has been previously seen in the literature.
5005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 000000.20.20.20.20.40.40.40.4 Train Dataset Size $p$-value ($\downarrow$) (a) Dataset Inference
10010010010015015015015020020020020025025025025030030030030000%$5$%$10$%$15$% ROUGE-L Length Extraction Rate ($\uparrow$) (b) Copyright Infringement
Best Single SetupComposite Prompt LengthsComposite Model SizesComposite Training Steps
Figure 6: (a) $p$-value for dataset inference (lower is better) across different dataset sizes. The results show the importance of exploiting prompt sensitivity and the significant improvement under the composite setup across different prompt lengths. (b) Extraction rate for different ROUGE-L length thresholds, marking potential copyright violations generated by the LLM. Extraction rates with composite setups are consistently higher than the single setup, highlighting the impact of multi-faceted access to the data.
6.2 Approximate Matching
As discussed in §3.2, evaluating extraction attacks under verbatim match can underestimate the true risk of extraction. To address this gap, we introduced approximate composite discoverable memorization (Definition 3.2), and will now analyze various similarity metrics $S$ to examine their behaviour under changing $\delta$, reported in Figure 5(c). Solely for this discussion, we increase the completion length $l_{x}=500$, to allow for meaningful extraction even with approximate matching.
Our results reveal intriguing trends. First, we analyze evaluations based on the Levenshtein ratio metric and observe that even the threshold of $\delta=0.95$ doubles the extraction attack rate compared to a verbatim match. This threshold signifies a minimum $95\%$ overlap between generated and original text. Even under such a strict threshold, the doubling of the extraction rate underscores the significant underestimation of extraction risks when relying solely on verbatim matches. As $\delta$ decreases, however, the extraction rate increases exponentially, as the Levenshtein ratio becomes less reliable under looser constraints. We also see similar trends for ROUGE-L scores.
Transitioning to other similarity metrics — longest common substring (LCS), Hamming distance, and $n$-gram matching — we find that even lower values of similarity ($\delta$) can contribute meaningfully to extraction attacks. We observe patterns of rising extraction rates similar to what we saw earlier with the Levenshtein distance. The diverse trends underscore the choice of approximation metric as highly context-dependent. A more thorough examination of which metrics best serve particular applications is left for future work.
6.3 Data Deduplication
A commonly recommended solution to extraction attacks and memorization is data deduplication, involving the removal of duplicate data entries within a dataset (Carlini et al., 2023). While costly, data deduplication represents a critical aspect of data curation and has been shown to mitigate extraction risks (Carlini et al., 2023). To understand the role of data deduplication in our discussion, we repeat our experiments using the models from Pythia trained on the deduplicated Pile dataset. The results are collected in Figure 5(b).
In line with existing literature, data deduplication reduces the extraction rate. Interestingly, however, we observe persistent trends: the presence of a stronger adversary due to multi-faceted dataset access. Thus, while beneficial, data deduplication does not alter our fundamental conclusions; real-world adversaries with multi-faceted access to the underlying data can extract substantial information even post-deduplication. Future work on incorporating more concrete frameworks like differential privacy is needed, to better understand such adversaries, particularly from the perspective of privacy protection under multi-access systems.
7 Case Studies with Stronger Adversary
We conclude by highlighting the value of our stronger adversary in various case studies.
7.1 Detecting Pre-Training Data
Extraction attacks are a primary tool in identifying whether certain data was included in a model’s training set. This can be valuable in assessing whether a model is trained on proprietary or sensitive data without permission, evaluating data contamination and leakage in various benchmarks, ensuring regulatory compliance to data governance policies, or even academic research to track the influence of datasets on the model.
While membership inference attacks (MIAs) have been commonly used to detect pre-training data, Maini et al. (2024) argues that MIAs are as good as random guessing when it comes to distinguishing between members and non-members from the same distribution. They show that these attacks learn how to distinguish between concepts, and not actual text, highlighting the importance of using IID data of members and non-members to appropriately perform dataset inference.
We borrow their setup and extend it to the composite setting by increasing the size of the training set for learning correlations. Thus, our composite setting can be alternatively seen as an augmentation technique for the training set. We record the $p$-value of the null hypothesis "the dataset was not used for training" for the Pile dataset in Figure 6(a), under different sizes of the original training data.
We find that the $p$-values for the composite setting with prompt lengths are noticeably lower than those for the baseline, especially at smaller dataset sizes. Thus, our adversary requires less data to achieve the same $p$-value as the baseline. The dataset inference setup by Maini et al. (2024) requires obtaining IID data that the data owners are certain was not used for training by the LLM, which can be difficult to find. Hence, reducing the amount of such data required can be extremely useful, which further emphasizes the value of considering a real-world adversary. Interestingly, we did not find similar strong trends for composite attacks across different model checkpoints. We believe this might be because membership inference information can change drastically across models, and thus combining information from multiple checkpoints does not help in learning better correlations.
7.2 Copyright Infringement
Copyright issues due to LLMs regurgitating their training data have been heavily studied in recent literature. Karamolegkou et al. (2023) discusses different thresholds for quoting a text ad verbatim that has been considered a violation of fair use, for example, $50$ words is a common threshold used for magazine articles, chapters, etc., while $300$ words is a common threshold used for books. The authors suggest using the longest common subsequence (ROUGE-L score length) as a measure of quantifying text reproduction and potential violations.
Following their reasoning, we record the distribution of ROUGE-L lengths for $2000$ randomly chosen examples in Figure 6(b), both for the strongest baseline as well as the composite settings. We find that a real-world adversary generates more potential copyright violations than the adversary in the literature, which highlights the underestimation of such risks. We note that copyright is a highly complex problem, and simply extracting data from the model might not necessarily constitute a copyright violation. However, our focus is on improving the technical underpinnings that are necessary for a fruitful discussion of copyright issues in LLMs.
7.3 PIIs Extraction Risk
Another commonly studied risk of memorizing training data is extracting personally identifiable information (PIIs). We use the setup of Li et al. (2024) to create our PII extraction test set from the Pile dataset. We use GLiNER (Zaratiana et al., 2024) to detect $2000$ unique PIIs in the Pile dataset, followed by cutting the sentence right before the PII to create the input prompt. These prompts were fed to the model, and the attack is considered successful if the correct PII is generated anywhere within the first 100 tokens, marking the risk of PII leakage (Li et al., 2024).
We record the extraction risk for the best single setup and composite extraction risks across model checkpoints and model sizes. Since the prompts in this setup are of varying lengths, we do not extend our changing prompt lengths setting to this case study. Similar to Definition 3.1, the composite PII extraction is considered successful if the PII is present in the generation of at least one of the models. The results are collected in the table below, and continuing previous trends, we see a noticeable increase in the extraction rate for an adversary with access to multiple checkpoints.
Setup Extraction Rate
Best Single Setup $22.16\%$
Composite Model Sizes $30.97\%$
Composite Training Steps $33.07\%$
8 Limitations and Future Work
By highlighting the multi-faceted access available to an adversary in the current LLM landscape, our work reveals a severe underestimation of information leakage risks in the existing literature. We emphasize the importance of explicitly considering the adversarial perspective and the composability of information leakage in extraction attacks.
Our real-world adversary is certainly more powerful but also more expensive than the adversary in the existing literature. Unlike our current setup, where we verify extraction using the ground truth, an adversary would need to justify both the cost of additional model generations and the expense of verifying extracted information. Therefore, future research should explore the cost-benefit trade-offs of multi-faceted access, focusing on when these added expenses may outweigh the benefits of new information extracted, particularly as we show diminishing gains with increased points of access.
As most of our analysis focuses on the risks posed by extraction attacks under the lens of discoverable memorization, future research should also explore how our findings translate to other forms of privacy attacks. Finally, our study addresses the threats posed by powerful real-world adversaries but does not propose specific defence methods. Further exploration is needed to navigate the current LLM ecosystem and mitigate the risks posed by these strong adversaries.
Acknowledgments
Funding support for project activities has been partially provided by Canada CIFAR AI Chair, and Google award. We also express our gratitude to Compute Canada for their support in providing facilities for our evaluations.